OpenStudy (jjuden):

Simplify completely:

OpenStudy (jjuden):

\[\sqrt{24x^{6}}\]

OpenStudy (anonymous):

sqrt of 24 is?

OpenStudy (jjuden):

\[2x ^{2}\sqrt{12x ^{2}};2x ^{3}\sqrt{6};4x \sqrt{6x ^{5}};2x \sqrt{6x ^{5}}\] answers

OpenStudy (anonymous):

x to power what ?